OpenStudy (anonymous):

Solve for all real values of x : (x^2 - 13x + 40)(x^2 - 13x + 42)/√(x^2 - 12x + 35)

OpenStudy (nikato):

is that a square root?

OpenStudy (mertsj):

How can you solve that? It is not an equation.

OpenStudy (anonymous):

yes

OpenStudy (nikato):

is it suppose to equal 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Mertsj its equal to 0 sorry

OpenStudy (nikato):

do you know how to factor?

OpenStudy (anonymous):

i did that but that is as far as i got

OpenStudy (nikato):

so what do you have after factoring?

OpenStudy (mertsj):

Consider the numerator. The only way a fraction can be 0 is if the numerator is 0 and the denominator is not 0. So set the numerator equal to 0. Then factor the numerator. Set each factor equal to 0. Solve each equation. Check to make sure that no answer causes the denominator to be 0.

OpenStudy (anonymous):

@nikato (x-8)(x-5)(x-7)(x-6)/√(x-7)(x-5)

OpenStudy (phi):

if you have \[\frac{ (x-8)(x-5)(x-7)(x-6)}{\sqrt{(x-7)(x-5)} }=0\] the top will be zero if any of the 4 factors is 0. for example if x-8=0 (which of course means x=8), the top will be 0 * other stuff= 0 so x=8 will be one answer *as long as the bottom is also not zero* as you can see, with x=8, the bottom is sqr(1*3) which is not zero. So x=8 is a solution try the others.

OpenStudy (anonymous):

so can i use 5 since it is x-5

OpenStudy (anonymous):

what numbers in the denominator will cause it to become undefined?

OpenStudy (phi):

yes, x=5 makes the top =0 but if the bottom is also 0 when x=5, you get 0/0 which is undefined, so not allowed

OpenStudy (anonymous):

how are you determining that? did you just plug in 5 for x??

OpenStudy (phi):

yes, just replace x with 5 and see what you get

OpenStudy (phi):

in case it is not clear, if you have (x-8)(x-5)(x-7)(x-6)=0 that can be broken down into 4 separate possibilities x-8=0 x-5=0 x-7=0 x-6=0 those are the four ways to make the top equal to zero. however, you have to reject any x value that ALSO causes the bottom to be zero because we do not allow 0/0 as an answer.

OpenStudy (anonymous):

so only 6 and 8 work right?

OpenStudy (phi):

yes, exactly

OpenStudy (anonymous):

are those all the zeros?